whats an equation with a variable on both sides that will give me the variable to be 24 ??
Answer by Edwin McCravy(20054)

This amounts to working a problem backward:

Start with the answer that you want to end up with:

           x = 24

Then multiply both sides by some arbitrary number,
say -2

         -2x = -48

Now add an arbitrary multiple of x, say 7x, to both sides:

    -2x + 7x = -48 + 7x

Combine the terms on the left:

          5x = -48 + 7x

That one will do, but if you like you can add an arbitrary
number to both sides, say 17

     5x + 17 = -48 + 7x + 17

Combine -48 and 17 on the right side: 

     5x + 17 = -31 + 7x


That one has the solution x = 24.  You can make as many equations
that have the solution x = 24 as you like, choosing different 
arbitrary numbers from the ones I chose.

Edwin